Cassini Significant Events -- 04/26/06 - 05/03/06

May 5, 2006

(Source: Cassini Project)


The most recent spacecraft telemetry was acquired Wednesday, May 3, from the
Goldstone tracking stations. The Cassini spacecraft is in an excellent state
of health and is operating normally. Information on the present position and
speed of the Cassini spacecraft may be found on the "Present Position" web
page located at http://saturn.jpl.nasa.gov/operations/present-position.cfm .

Wednesday, April 26 (DOY 116):

Unraveling the Twists and Turns of Saturn's Rings was the topic at the
Cassini-Huygens Analysis and Results of the Mission (CHARM) teleconference
on April 24. The presentation included a rings tutorial and recent
Ultraviolet Imaging Spectrograph (UVIS) results.

Orbit trim maneuver (OTM) # 58 was performed today. This is the approach
maneuver setting up for the Titan 13 encounter on April 30. The reaction
control subsystem burn began at 8:14 AM PDT. A "quick look" immediately
after the maneuver showed the burn duration was 53.1 seconds, giving a
delta-V of approximately 77.4 mm/s. Telecommunications reported a momentary
1dB drop in signal strength during the roll turn. They are currently
investigating. All other subsystems reported nominal performance after the
maneuver.

Science activities today centered on the RADAR obtaining distant Titan
radiometer science and calibration data over a range of phase angles. This
observation is designed to reveal any phase variations that may be present
in Titan radiometry and scatterometry data.

The Science Planning leads for S22 briefed Program Management on the
contents of the S22 sequence, and any issues that remain open that will be
transferred into the next phase of sequence development. A Project Briefing
is the final step in the Science Operations Plan update process.

As was reported on Monday, April 24, the Program Scientist was to make a
final decision on what data to protect from being overwritten in the event
of a problem with the DSN after the Titan 13 (T13) flyby. He has asked
Science Planning, the S20 sequence leads, and members of the Spacecraft
Operations Office if it might be possible to preserve at least one block of
the unique F ring data in the post T13 time period if such an anomaly were
to occur. It was decided that the easiest strategy to implement was to save
all of the Visual and Infrared Mapping Spectrometer (VIMS) data on the day
after the T13 flyby, and to zero out all of the other instruments if a
second playback is required. Since the total VIMS data volume is relatively
small, the VIMS data, along with the Synthetic Aperture RADAR (SAR)/UVIS
block, could all be played back over the next downlink. This strategy will
preserve the key science data for both Titan and VIMS rings.

In the event of a problem, a file will be uplinked that will 1) create a new
data policing table so no science is recorded for remainder of the downlink
on DOY 121, 2) enable this table immediately once the file reaches the
spacecraft, and 3) modify Table 19 for the following observation period on
DOY 121/21:14 - 122/19:43 SCET. Only VIMS packets will be recorded.

A second file to be uplinked sometime around the start of track on DOY
122/21:10 UTC will move the pointers to recover the lost RADAR/UVIS data.
The strategy behind this is not straightforward and does come with a risk.
Anytime the pointers are moved there is a chance something can go wrong and
data unintentionally lost. The CDS team met to discuss this and reduce the
risk. Once the pointers are moved, all science data from then on will be
played back as normal.

Since all of this is a contingency plan, it would be nice if there were no
problems after all, and the files not uplinked. Only time will tell, but
the files will be approved and ready to go just in case.

Thursday, April 27 (DOY 117):

It was reported on Tuesday, April 18, that the S21 leads were planning to
schedule a test in the Integrated Test Laboratory (ITL) in April to ensure
that de-registration of S20 part 2, and the registration of the S21
background sequence occur as expected. It is desired to use the latest
version of the S21 sequence for this "loading test". Based on the
development schedule, the preliminary #2 merge will be available on May 22.
If development continues on its present course, this will be the final merge
in the sequence development process as a final phase will be unnecessary.
ITL staff are currently reviewing the test schedule to determine if the week
of May 22 is feasible for the test.

Friday, April 28 (DOY 118):

An ACS Reaction Wheel Assembly (RWA) exercise of the backup wheel was
performed today. In this activity, RWA-3 is commanded to +100 rpm, -100
rpm, 0 rpm, then turned off.

The Cosmic Dust Analyzer (CDA) instrument team has delivered their new
flight software to the Project Software library. The Software Requirements
and Certification Review (SRCR) is scheduled for Wednesday, May 3.

Sunday, April 30, (DOY 120):

At 2006-120T20:58 spacecraft event time, Cassini flew past Titan at an
altitude of 1855 km for the 14th targeted Titan encounter of the tour.
During this event, the Imaging Science Subsystem observed particle
properties, vertical distributions, wind and cloud motions, and monitored
lightning and aurora activity, the Composite Infrared Spectrometer (CIRS)
obtained information on trace constituents in Titan's stratosphere by
integrating on the limb at two positions, the Ultraviolet Imaging
Spectrograph (UVIS) observed the star Beta Ori as it was occulted by Titan's
atmosphere, and the Visible and Infrared Mapping Spectrometer (VIMS)
obtained new high-resolution images that will help understand Titan's
geology and the fate of CH4.

The entire suite of Magnetospheric and Plasma Science (MAPS) instruments
performed observations in the immediate vicinity of Titan studying
magnetospheric and ionospheric interactions.

Finally, RADAR collected altimetry, radiometry, and scatterometry data, and
performed low and high resolution Synthetic Aperture RADAR (SAR) imaging of
Titan's surface. The SAR swath cuts right across Xanadu, as well as some
areas where there exists good ISS and VIMS coverage, enabling useful
comparative studies between disciplines.

Monday, May 1 (DOY 121):

Remember those contingency plans discussed earlier in this report?
Wellllllllll.

Monday was a busy day. Cassini had a DSN pass over Goldstone on Saturday,
nothing on Sunday as planned, and our next tracks were the Madrid array on
Monday, May 1. What follows is a mini diary of a spacecraft anomaly. The
following has been extracted from various e-mails jetting through the
project today. Read on:

DIARY START

Sequence Lead:

The planned downlink from the spacecraft (S/C) did not go as expected this
morning. From the beginning of the track, we have been unable to see any
signal from the orbiter. The uplink was swept multiple times, and a signal
from the low-gain antenna has been looked for with no effect. I am callling
an anomaly meeting for this morning at 9:00 PDT.

Spacecraft Operations Office (SCO):

A loss of downlink anomaly occurred on May 1, 2006. No one-way data was
received at the beginning of track at around 6:30 a.m. PDT. We started the
uplink drive on time at 6:40 a.m., but stopped after 50 minutes when the
no-downlink anomaly continued. After looking for all combinations of
spacecraft safing with DSS 63, DSS 55, and Radio Science Open Loop
receivers, SCO decided to wait for nominal lockup at the two-way time.

Sequence Lead:

Good news! At the time we were scheduled to go 2-way with the S/C we
regained downlink. Data is now coming in as expected. Telemetry indicates
there was a Solid State Power Switch (SSPS) trip on the unit that powers the
Ultra Stable Oscillator (USO), and the USO is currently powered off. This
means that a 2-way light time after the transmitter was turned back off,
during the recovery efforts, we will go back to 1-way, and thus will lose
the downlink again, for approximately 2 hours. Since this will almost
definitely impact the SAR data, we will be sending up the contingency
command to zero out data policing during this pass, and zero out everything
except VIMS during the following observation period. The SAR data will then
be played back during tomorrow's pass.

SCO is currently planning to investigate whether the SSPS trip was merely a
cosmic ray hit or whether there is something actually wrong with the USO,
and we should know that by the end of the current pass. I will forward that
info as soon as I have it. We are trying to get our DSS-15 track tomorrow
extended to our rise time, so that we can go 2-way as early as possible, in
case the USO is damaged.

Spacecraft Operations Office:

It turns out that the SSPS for the USO was tripped, causing no one-way
downlink carrier or data. This tripped switch condition is consistent with
ones seen in the past. This is the sixteenth trip seen to date, the third
trip of a switch that was ON at the time, and the second trip this year. The
previous trip occurred very recently on March 2, 2006. They are predicted
to occur at a rate of about two per year, and are most likely caused by
Galactic Cosmic Rays.

Commands were built and uplinked to cycle the SSPS OFF to clear the trip,
then commanded it back ON. The S/C is operating nominally following this
activity.

Sequence Lead:

The file to zero out data policing during the current pass and for all but
VIMS on the next observation period has been sent with a bit-1 time of
121T17:34:55. We will confirm that this file reached the spacecraft at
approximately 20:00.

The DSN was able to extend tomorrow's track for Cassini, starting it at
station rise at 19:15 rather than at our allocated time of 21:10. This will
allow us to do an early uplink, and go 2-way sooner than we would otherwise,
which will be a boon, if cycling the SSPS does not clear the issue up.

Instrument teams have been reporting nominal operations for their
instruments. All are awaiting playbacks to determine what data was
overwritten.

I will keep you posted on any further development.

Tuesday, May 2 (DOY 122):

DIARY CONTINUES

Spacecraft Operations Office:
Telemetry indicates that the command file to move the pointers to replay
Titan data has executed. The pointers were moved to the appropriate
locations discussed in yesterday's meeting with CDS, Science Planning,
Mission Planning, Uplink Operations, and RADAR. Now we just wait for the
data ...

RADAR Instrument Team:

Just received the first RADAR packet on the ground. The time lines up at
about -12 minutes from closest approach. So looks like the first ~90 Mbit
got overwritten as predicted. Hopefully the end point is correct. Looks
good and as expected so far!!

UVIS Instrument Team:

The UVIS T13 occultation data is on the ground. The data looks good except
for an unfortunate 1-minute data gap just when the signal starts to
decrease. It is assumed that this is due to the S/C going 2-way or to a
telemetry rate change. In any case nice recovery work! Thanks to all
involved in this data recovery.

RADAR Instrument Team:

Great pointer work!!

The end point was also correct!

We got all our data from about -12 minutes up to where we receive data
yesterday. There was a small overlap, so there are no gaps in the data that
we do have. The only missing data is the -20 to -12 minutes of SAR that was
overwritten. So basically ~600 out of the expected 700 Mbit was recovered.

Great work! Thanks is probably not enough.........but thanks!

END DIARY

The final sequence development process for S22 kicked off today. The
process will run for approximately 11 weeks and will begin execution on July
17.

Cassini Outreach has reported the release of the DVD "Ring World 2". The DVD
includes both English and Spanish language versions, as well as closed
captioning.

Copies have been shipped out to members of the International Planetarium
Society, the Navigator Program's Night Sky Network, NASA Solar System
Ambassadors, Space Place's museum network, the USGS, and throughout the
Cassini mission to flight team members who give Cassini outreach
presentations. A RW2 version for video Ipod is now available and may be
downloaded from:
http://saturn.jpl.nasa.gov/multimedia/videos/video-details.cfm?videoID=114

Wednesday, May 3 (DOY 123):

OTM #59 was performed today. This is the cleanup maneuver from the Titan13
encounter on April 30. The main engine burn began at 7:45 PM PDT. A "quick
look" immediately after the maneuver showed the burn duration was 3 seconds,
giving a delta-V of approximately 0.47 m/s. All subsystems reported nominal
performance after the maneuver.

At an SRCR held today, the new CDA flight software was approved for
operations. The planned uplink date is June 5, with on-board checkout
scheduled to occur on DOY 165, 9 days later.

A beautiful color picture of Saturn is Astronomy Picture of the Day today.

Upcoming Events:

On May 6, over 70 Saturn Observation campaign members in 33 states and 10
countries are conducting daytime solar viewing activities and/or nighttime
stargazing to celebrate International Astronomy Day. These events range
from small neighborhood planet viewings to large events. Weather permitting,
Saturn will be a primary viewing target. Contact Cassini Outreach for
specifics, or contact your local astronomy club, science center, or college
to see if there is an event in your community.

JPL Open House - May 20-21, 2006: Expected Attendance: 50,000
Cassini will have a 30'x50' exhibit booth near the main entrance of JPL. The
booth will highlight mission science results through graphic displays and
multimedia presentations. The both will be staffed with project employees
who will be available to answer questions. A kids' area will feature
"temporary tattoos" of Saturn. Interactive features include "Take your photo
on Titan" with a full scale Huygens probe model placed in a simulated
landing environment as well as "spacecraft tours" using the 1/2-scale
Cassini model. A "Moon Walk" will guide Open House visitors past highlights
of Saturn's moons. The moonwalk will link the booth and the 167 conference
room where Ring World will be shown.